# The Speech-LLM Takes It All: A Truly Fully End-to-End Spoken Dialogue State Tracking Approach

### v1

@@ 4 Conclusion
In this paper, we have proposed a fully E2E approach to Spoken Dialog State Tracking, drawing inspiration from Speech-LLMs. In contrast to traditional multimodal context approaches, we show that it is possible to use the entire spoken conversation as input (until the current turn) and achieve state-of-the-art results. We also have performed a fine-grained analysis to illustrate the causes of improvements brought by using a full spoken context: less error propagation through the dialog and better performance on the most challenging slots. In future work, a more sophisticated and compact handling of the spoken context may be explored. Moreover, scaling the used model would be a promising extension.

### v2

@@ 4.  Conclusion
In this paper, we have proposed a fully E2E approach to Spoken Dialog State Tracking, drawing inspiration from Speech-LLMs. In contrast to traditional multimodal context approaches, we show that it is possible to use the entire spoken conversation as input (until the current turn) and achieve state-of-the-art results. We also performed a fine-grained analysis to illustrate the causes of improvements brought by using a full spoken context: less error propagation through the dialog and better performance on the most challenging slots.
At the same time, using the full spoken history entails practical trade-offs. Most notably, the increased memory usage and latency motivate the design of more efficient context representations. In future work, a more sophisticated and compact handling of the spoken context may be explored. Moreover, scaling the used model would be a promising extension. We further plan to investigate streaming inference and integration within production pipelines. We also are interested in enlarging the scope of the evaluation beyond SpokenWOZ, as additional human-spoken DST benchmarks may become available. Finally, we aim to release code and trained models to facilitate reproducibility.
5.  Limitations
While our experiments show that providing the full spoken history is beneficial for spoken DST, some limitations need to be mentioned. First, the memory and compute footprint of the full-spoken setting grows with dialog length. Indeed, each additional spoken turn expands the input sequence to process. It then increases attention cost (quadratically) and thus end-to-end latency. This problem is especially noticeable on long dialogs or when having production-level constraints on hardware (GPU memory can become a bottleneck even with relatively small context windows). We can note that our proposed compression variant helps tackle this issue, yet it may imply the loss of some fine-grained cues, such as named entities, numbers, or accurate time slots).
Then our models are effective enough for an offline research environment. Yet, they remain far from a delivery-level. Real-world TOD systems often require to enable streaming (e.g. partial hypotheses, incremental state updates) with small latency. The scope of the paper was not including the end-to-end latency and the failure modes under streaming constraints, which are all prerequisites for reliable deployment and may be included in future work.
Finally, our empirical analysis is focused on SpokenWOZ. To the best of our knowledge, it is the main benchmark built from human conversational speech for DST. We may extend the study to other benchmarks such as Spoken MultiWOZ, but it would require care in interpretation. This kind of corpora is mainly composed of synthesized voice for training and validation. Human speech is only in evaluation splits. So a large part of the dataset contains acoustic characteristics that are very different from fully human-spoken data. Thus, performance measured on such datasets may reflect robustness to TTS artifacts as much as to real conversational variability.
Future works may help to address these three limitations. We may explore more memory-efficient context representations (such as hierarchical or selective retention), dig on streaming compatible inference, and enlarge the evaluation scope to synthesized corpora with great care for acoustic quality.